OpenStudy (anonymous):

The hypotenuse of a 30°- 60°- 90° right triangle measures 22 cm. What are the lengths of the longer leg and shorter leg? @tcarroll010

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

short leg =11 longer leg= 11\[\sqrt{3}\]

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

The shorter leg of a 3-60-90 degree right triangle is half the hypotenuse, so the shorter leg is 11. Let "x" be the length of the longer leg, then: x^2 + 11^2 = 22^2 x^2 = 22^2 - 11^2

OpenStudy (anonymous):

you multiply the 11 and the square root of three together to get longer leg

OpenStudy (anonymous):

x^2 = 22^2 - 11^2 gives u 19.05?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so thats the longer leg?

OpenStudy (anonymous):

yes.